### v1

# QuMAB: Query-based Multi-annotator Behavior Pattern Learning

Abstract.
Multi-annotator learning traditionally aggregates diverse annotations to approximate a single “ground truth”, treating disagreements as noise. However, this paradigm faces fundamental challenges: subjective tasks often lack absolute ground truth, and sparse annotation coverage makes aggregation statistically unreliable. We introduce a paradigm shift from sample-wise aggregation to annotator-wise behavior modeling. By treating annotator disagreements as valuable information rather than noise, modeling annotator-specific behavior patterns can reconstruct unlabeled data to reduce annotation cost, enhance aggregation reliability, and explain annotator decision behavior. To this end, we propose QuMATL (Query-based Multi-Annotator Behavior Pattern Learning), which uses lightweight queries to model individual annotators while capturing inter-annotator correlations as implicit regularization, preventing overfitting to sparse individual data while maintaining individualization and improving generalization, with a visualization of annotator focus regions offering an explainable analysis of behavior understanding. We contribute two large-scale datasets with dense per-annotator labels: STREET (4,300 labels/annotator) and AMER (average 3,118 labels/annotator), the first multimodal multi-annotator dataset.
Multi-annotator learning, Annotator tendency, Behavior patterns, Multi-annotator datasets
ccs: Computing methodologies Multi-task learning
1. Introduction
Figure 1. Paradigm shift from sample-wise aggregation to annotator-wise behavior modeling. (a): Sparse annotation matrix showing each annotator labels a small subset of samples with disjoint coverage. (b): Traditional sample-wise aggregation makes a questionable single “ground truth” prediction, potentially losing individual information to limit behavior analysis. (c): Our annotator-wise behavior modeling captures each annotator’s behavior patterns longitudinally across their labeled samples, enhancing reliability of aggregation via reconstructed unlabeled data in annotation matrix, offering explainable analysis of behavior understanding.
In real-world multi-annotation scenarios, such as medical image analysis (Liao et al., 2024), sentiment analysis (Lian et al., 2023), and visual perception (Zhang et al., 2023e), different annotators often provide different labels to the same sample (Raykar et al., 2010) due to different personal backgrounds, subjective interpretations, and preferences. Traditional multi-annotator learning focuses on learning different characteristics (e.g., confusion mode (Tanno et al., 2019a), agreement (Wang et al., 2023), expertise level (Ji et al., 2021a)) from multiple annotators, then treating these discrepancies as bias or noise to elimant for achieving aggregation to approximate a single “ground truth” prediction (Yan et al., 2014; Jinadu et al., 2023).
However, the reliability of this paradigm faces two fundamental challenges: (1) In subjective domains such as emotional or impression assessment, there often exists no absolute ground truth—making this aggregation itself questionable (Raykar et al., 2010; Snow et al., 2008). (2) In real-world crowdsourcing, each annotator labels only a small fraction of the data, with most samples receiving annotations from different, often disjoint annotator subsets. This sparse and fragmented coverage makes aggregation statistically unreliable, as there is insufficient overlap to establish robust consensus patterns (Nørregaard and Derczynski, 2022).
Therefore, we argue for a shift in focus, i.e., from sample-wise to annotator-wise (Figure 1): instead of treating sample-wise annotator disagreements as noise to be averaged away, we propose modeling individual annotator behavior patterns as annotator-wise valuable information. These patterns capture consistent differences in judgment arising from expertise, preference, or perspective. By longitudinally learning reliable annotator-specific models—tracking each individual’s behavior patterns across consecutive sample-label pairs per annotator rather than aggregating multiple annotators per sample—we unlock three compelling advantages: (1) Cost reduction: reconstructing unlabeled data enables comprehensive annotation coverage; (2) Enhanced reliability: aggregating over reconstructed sufficient coverage of sample-label pairs yields statistically more robust consensus than sparse, fragmented annotations; (3) Behavioral insights: understanding the behavior patterns underlying each annotator’s decisions and explaining sources of disagreement.
Currently, existing research focusing on multi-annotator behavior pattern modeling and presenting explainable analysis of behavior understanding is sparse. Some works similar to it have attempted to model individual annotators through various techniques (e.g., MaDL’s confusion matrices (Herde et al., 2023) or PADL’s Gaussian distribution fitting (Liao et al., 2024)) to understand more about individual annotator patterns. However, their aggregation-oriented mechanism (e.g., PADL’s meta-learning, and MaDL’s jointly optimizing consensus and annotator classifiers) may average annotator perspectives to lose annotator-specific information, influencing individual behavior modeling. Otherwise, if completely independent modeling individual annotator, it preserves sufficient individual information but easily suffers from overfitting during the annotator model is trained on its small set of labels as described previously. Moreover, existing models (Chen et al., 2019) lack explainable analysis of behavior understanding, or they only implicitly reveal trends where certain annotators play a larger role in the prediction (Cheng et al., 2023).
Our key insight is that annotators, despite their individual differences, often share behavioral structures. By capturing inter-annotator correlations, we can leverage the collective patterns as implicit regularization, constraining individual models from overfitting while preserving unique characteristics. To this end, we propose a novel query-based architecture QuMATL (Query-based Multi-Annotator Behavior Pattern Learning), hypothesize that annotator judgment differences arise from their varying degrees of focus on different regions of the input content (e.g., focusing on different image patches). Each annotator is represented by a learnable query that interacts with input features via cross-attention to effectively model individual behavior patterns. Lightweight query significantly reduces computational cost compared to separate conventional models. Crucially, all annotator queries also interact through shared self-attention to capture inter-annotator correlations as a form of implicit structural regularization. This constrains inter-annotator representations to follow similarity patterns derived from annotations, preventing individual representations from drifting too far from the group and promoting mutual enhancement. This mechanism prevents overfitting to sparse individual data while maintaining individualization, improving generalization and robustness in individual annotator modeling, particularly under sparse annotations. Additionally, the cross-attention weights provide a visualization of annotator focus regions, offering an explainable analysis of behavior understanding.
Furthermore, we contribute two new large-scale datasets with dense per-annotator labels: STREET (city impression assessment, 4,300 labels/annotator) and AMER (video emotion recognition, average 3,118 labels/annotator). These datasets provide a high-value longitudinal annotation perspective for understanding and evaluating individual annotator behavior patterns, offering valuable data to the community and further researchers. It is worth noting that AMER is the first multi-annotator multimodal dataset in this field. Our work makes the following contributions:
  • A paradigm focus shift in multi-annotator learning: We introduce a paradigm shift from sample-wise consensus aggregation to annotator-wise behavior modeling. By treating annotator disagreements as valuable information rather than noise, modeling annotator-specific behavior patterns can reconstruct unlabeled data to reduce annotation cost, enhance aggregation reliability, and explain annotator behavior.
  • A novel query-based architecture: We propose QuMATL, which uses lightweight queries to model individual annotators while capturing inter-annotator correlations as implicit regularization, preventing overfitting to sparse individual data while maintaining individualization and improving generalization, with a visualization of annotator focus regions offering an explainable analysis of behavior understanding.
  • Two new large-scale datasets: We contribute STREET (4,300 labels per annotator) and AMER (average 3,118 labels per annotator) datasets with denser per-annotator labels than existing resources, offering a longitudinal perspective for understanding individual annotator behaviors. AMER is the first multimodal multi-annotator dataset.
Dataset Dataset description Modality # samples per annotator
QUBIQ-kidney (Menze et al., 2020) kidney image image 24
QUBIQ-tumor (Menze et al., 2020) brain tumor image image 32
QUBIQ-growth (Menze et al., 2020) brain growth image image 39
QUBIQ-prostate (Menze et al., 2020) prostate image image 55
CIFAR-10H (Peterson et al., 2019b) object recognition image 200
MUSIC (Rodrigues et al., 2013a) music genre classification audio 2$\sim$368
MURA (Rajpurkar et al., 2017) radiographic image image 556
RIGA (Almazroa et al., 2017) retinal cup and disc segmentation image 750
LIDC-IDRI (Armato III et al., 2011) lung nodule image image 1,018
STREET (Ours) city impression evaluation image 4,300
AMER (Ours) video emotion recognition audio, video, text 970$\sim$5,202
Table 1. Dataset comparison. Compared to existing datasets, our datasets contain a greater number of samples annotated by each annotator, helping promote multi-annotator behavior pattern modeling. AMER is the first multimodal multi-annotator dataset.
2. Related Work
2.1. Multi-annotator Behavior Modeling Paradigm
To the best of our knowledge, the multi-annotator behavior modeling paradigm problem has not yet been investigated. Traditional multi-annotator learning focuses on estimating consensus or ground-truth labels from multiple noisy annotations. These include early probabilistic models (Dawid and Skene, 1979), EM algorithms (Whitehill et al., 2009), Gaussian models (Rodrigues et al., 2014), and biased estimation (Welinder et al., 2010). Tanno et al. (Tanno et al., 2019b) proposed modeling annotator confusion matrices as learnable parameters in neural networks. Cao et al. (Cao et al., 2019) introduced max-MIG to learn from multiple annotators. NEAL (Chen et al., 2023) employs neural expectation-maximization to jointly learn annotator expertise and true labels. Later methods used probabilistic frameworks to aggregate multiple annotations into a consensus or ground-truth label by confusion matrix (Tanno et al., 2019a), agreement distribution (Wang et al., 2023), and Gaussian distributions (Liao et al., 2024). This sample-wise aggregation paradigm often treats annotator disagreements as noise to be averaged away rather than valuable information (Yan et al., 2014; Jinadu et al., 2023). In contrast, our introduced annotator-wise modeling paradigm treats annotator disagreements as valuable information for modeling annotator-specific behavior patterns, enhancing aggregation reliability, and explaining annotator behavior.
2.2. Multi-annotator Behavior Modeling Architecture
Previous studies have attempted to model individual annotators through various techniques: D-LEMA (Mirikharaji et al., 2021) trains annotator models on non-contradictory subsets with spatial weights for noise handling; PADL (Liao et al., 2024) models annotator preferences via Gaussian assumptions in its Human Preference Module (HPM) and employs a Sample Embedding Module (SEM) for meta-classification; MaDL (Herde et al., 2023) jointly optimizes ground truth classifiers and annotator models through weighted embeddings. However, their aggregation-oriented mechanisms compromise individual behavior modeling by averaging annotator perspectives: D-LEMA sacrifices annotator-specific patterns for fusion objectives, PADL forces individual distributions to converge at the cost of behavioral details, and MaDL smooths individual characteristics to minimize consensus loss. Meanwhile, existing efforts on explainable analysis of annotator behavior understanding remain limited. Some works provide insights, e.g., TAX (Cheng et al., 2023) associates convolutional kernels with prototype libraries for pixel-level annotation decisions, MAGI (Zhang et al., 2023a) leverages annotator explanations to address noisy annotations, and Schaekermann et al. (Schaekermann et al., 2019) analyze factors contributing to disagreements. However, they only reveal annotators’ trends in aggregation or analyze isolated factors without behavioral analysis. In comparison, our architecture models individual annotators via lightweight queries, leveraging inter-annotator correlations as regularization against overfitting while preserving individualization, with attention visualization analyzing behavioral patterns.
2.3. Multi-annotator Datasets
Most existing multi-annotator datasets often have only a small subset of samples with consecutive annotations from consistent annotator IDs. CIFAR-10H (Peterson et al., 2019a), based on the CIFAR-10 (Krizhevsky, 2009) dataset, includes 10,000 test samples labeled by 2,571 annotators, but each annotator ID has on average only about 200 consecutive labels. LabelMe (Rodrigues and Pereira, 2018) includes an average of approximately 42.4 consecutive labels per annotator ID. Audio dataset Music (Rodrigues et al., 2013b) contains an average of about 46.1 consecutive labels. The medical datasets are commonly used in multi-annotator studies, and consecutive annotator labels are even sparser. QUBIQ (Ji et al., 2021b), a dataset for quantifying uncertainty in biomedical image segmentation, includes four distinct segmentation datasets with an average of only 40 samples, and even then, annotator IDs have only around 8 consecutive labeled samples each. For longitudinal annotator behavior understanding, we contribute two new large-scale datasets with dense per-annotator labels: STREET (city impression assessment, 4,300 labels/annotator) and AMER (video emotion recognition, average 3,118 labels/annotator). AMER is the first multimodal multi-annotator dataset.
Figure 2. QuMATL for image. A frozen pre-trained image encoder extracts features, which interact with annotator-specific learnable queries in the Image Q-Former through cross-attention, producing annotator-specific features for classification. Different annotators’ cross-attention weights reflect the differences in the image patches they focus on and support visual interpretation.
3. Dataset Construction
We contribute two new large-scale datasets [^1] We are negotiating to publish both datasets after acceptance.: STREET (city impression assessment) and AMER (multi-modal emotion recognition) in this paper. Table 1 compares the current multi-annotator datasets. We observe that in existing datasets, the number of samples annotated by each annotator is relatively small, and there is a lack of multi-annotator multimodal datasets. For example, in the RIGA dataset (Almazroa et al., 2017), each annotator labels 750 samples, while in the CIFAR-10H dataset (Peterson et al., 2019b), each annotator labels 200 samples.
(1) STREET is an urban perception dataset with multi-annotators, which contains $4,300$ high-resolution images covering various urban elements, such as streets, public spaces, and infrastructure. The images were captured during a series of city strolling surveys, which aim to analyze emotions in relation to various factors associated with the city. The surveys were conducted by an organization to which one of our co-authors belongs. Voluntary participants walked around their own familiar city, took photos of various factors that may affect their subjective feelings (i.e., happiness/health), and assigned labels related to these feelings to each image (though we do not use these labels but ones assigned by crowd workers in our experiments). Thirteen survey sessions were conducted in five different cities (three urban areas and two suburban areas). A total of 327 participants, ranging in age from their 10s to 60s, took part in the survey. Each session lasted about one hour, with each participant taking an average of 12.6 photos. We outsourced the annotation process to a company, which selected $10$ annotators with balanced age, gender, and location diversity on a platform similar to Amazon Mechanical Turk, assessing five perception dimensions: happiness, healthiness, safety, liveliness, and orderliness, using a $6$-point scale ($-3$ to $+3$). Each annotator spent approximately three weeks on their annotations. This multi-annotator dataset provides comprehensive human perception data for urban environments, enabling the quantitative analysis of environmental features and their emotional impact.
(2) AMER is a multimodal emotion dataset; the raw data is sourced from MER2024 (Lian et al., 2024), which contains $5,207$ video samples from movies and TV series, with multi-annotator emotion labels. Each sample typically contains one person, with relatively complete speech content. We annotated AMER using the open-source software Label Studio (Tkachenko et al., 2025). We hired $15$ annotators, who were students of our co-author’s institution, and underwent a training session with $10$ samples. We retained $13$ annotators after screening out careless and irresponsible ones. Each annotator completed the task in approximately two weeks, with scheduled breaks to maintain annotation quality, where each annotator selects the most likely label from 8 candidate labels, i.e., worry, happiness, neutrality, anger, surprise, sadness, other, and unknown. Among all annotators, $10$ annotators show consistent participation, each providing approximately $970$ to $1,096$ labels, while the remaining 3 annotators contribute over $5,000$ labels each. This rich multi-annotator setup provides reliable emotion annotation results and allows for a robust evaluation of emotion recognition performance.
4. Methodology
We propose QuMATL, a query-based architecture designed to model the behavior patterns of individual annotators. We take the image classification task to illustrate our image-specific architecture for image inputs from the STREET dataset, which consists of a frozen image encoder, annotator-specific learnable queries, an image Q-Former, and annotator-wise classifiers, as shown in Figure 2. For video inputs from the AMER dataset, we describe a video-specific architecture in the supplementary material.
Given an image input $I\in\mathbb{R}^{H\times W\times 3}$, a frozen pre-trained image encoder (Sun et al., 2023) first extracts image features, which are then fed into the Image Q-Former (Li et al., 2023). For each annotator $A_{k}$ ($k=1,\dots,n$), we assign a learnable query token for modeling. These queries are first processed by a shared self-attention layer to allow them interact with each other to implicitly capture inter-annotator correlations, and then interact with image features through multi-head cross-attention (typically with 12 heads), enabling each query to access diverse attention perspectives and produce individualized representations that reflect each annotator’s potential focus and decision process. The resulting representations are mapped through a fully connected layer and passed to each annotator’s classifier for prediction.
This query-based design is motivated by the hypothesis that annotator judgment differences arise from their varying degrees of focus on different regions of the input content (e.g., focusing on different image patches). Through learnable queries and cross-attention, our model effectively captures these individualized behavior patterns. Additionally, representing each annotator with a lightweight query significantly reduces computational cost compared to separate conventional models.
Meanwhile, the mechanism of capturing inter-annotator correlations acts as a form of implicit structural regularization. This constrains inter-annotator representations to follow similarity patterns derived from annotations, preventing individual representations from drifting too far from the group and promoting mutual enhancement. It also prevents overfitting to annotator-specific noise while preserving individual differences of behavior patterns. As a result, the model achieves improved generalization and robustness in individual annotator modeling, particularly under sparse annotations.
For qualitative understanding, we visualize the cross-attention weights from the Image Q-Former to reveal annotator-specific focus regions. These weights indicate which image patches different annotators focus on when making predictions. Figure 2 illustrates results on the STREET dataset (city impression classification), annotators exhibit distinct spatial focus: annotator n attends more strongly to the two people holding hands in the center. This contrast reflects how annotators may interpret emotional cues differently based on their focus: annotator n assigns a higher score to the “happiness” dimension compared to annotator 1, suggesting that variations in focus on semantically positive regions may contribute to their differing judgments.
4.1. Loss Function
Finally, as shown in Figure 2, the total training loss $\mathcal{L}_{\text{total}}$ for the proposed multi-annotator classification model, QuMATL, is defined as the sum of individual cross-entropy losses for each annotator:
(1) $\mathcal{L}_{\text{total}}=\sum_{k=1}^{n}\mathcal{L}(\hat{y}_{k},y_{k}),$
where each annotator $A_{k}$ has a specific predicted probabilities $\hat{y}_{k}\in[0,1]^{C}$, a reference label $y_{k}\in\{0,1\}^{C}$ in one-hot vector representation, and $C$ is the number of classes.
5. Experiment
We conduct extensive experiments to evaluate our QuMATL, including modeling individual annotators’ behavior patterns, assessing their utility for consensus prediction, testing applicability under sparse annotations, and complementing with qualitative visualization analysis. We compare against three representative baselines: D-LEMA (Mirikharaji et al., 2021), an ensemble-based multi-annotator learner; PADL (Liao et al., 2024), which fits Gaussian distributions for each annotator; and MaDL (Herde et al., 2023), which models annotator-specific confusion matrices. To ensure our model captures annotator-specific patterns rather than shared encoder features, we also include a Base variant with only the encoder and classifiers. Results are reported on two novel datasets. Accuracy and $F_{1}$ score (Zhan et al., 2019) are used as evaluation metrics. Note that additional experiments, including model efficiency, a faithfulness-oriented interpretability analysis, extended results, and further discussion, are provided in the supplementary material.
Metric Methods $A_{1}$ $A_{2}$ $A_{3}$ $A_{4}$ $A_{5}$ $A_{6}$ $A_{7}$ $A_{8}$ $A_{9}$ $A_{10}$ $A_{11}$ $A_{12}$ $A_{13}$ Avg CoPr
ACC Base 0.30 0.29 0.43 0.25 0.24 0.20 0.26 0.19 0.34 0.10 0.41 0.48 0.19 0.28 0.35
D-LEMA 0.86 0.88 0.85 0.87 0.89 0.86 0.88 0.87 0.85 0.86 0.45 0.51 0.33 0.78 0.55
PADL 0.89 0.90 0.88 0.93 0.87 0.91 0.86 0.94 0.89 0.88 0.47 0.54 0.35 0.79 0.52
MaDL 0.93 0.91 0.90 0.89 0.90 0.88 0.90 0.89 0.87 0.92 0.50 0.53 0.37 0.80 0.57
Ours 0.94 0.93 0.93 0.94 0.94 0.92 0.93 0.95 0.93 0.93 0.59 0.61 0.40 0.84 0.60
$F_{1}$ Base 0.26 0.27 0.40 0.22 0.23 0.17 0.24 0.18 0.31 0.08 0.35 0.41 0.14 0.25 0.32
D-LEMA 0.84 0.87 0.81 0.84 0.86 0.85 0.86 0.82 0.83 0.84 0.38 0.44 0.27 0.73 0.52
PADL 0.86 0.88 0.85 0.91 0.83 0.89 0.82 0.92 0.85 0.86 0.41 0.50 0.29 0.76 0.49
MaDL 0.90 0.85 0.87 0.86 0.87 0.85 0.87 0.86 0.84 0.92 0.45 0.48 0.33 0.77 0.54
Ours 0.91 0.91 0.90 0.92 0.91 0.90 0.89 0.93 0.91 0.93 0.54 0.55 0.34 0.81 0.57
Table 2. The accuracy (ACC) and $F_{1}$ score evaluate results on the AMER dataset. We assess performance for individual annotator modeling (each annotator $A_{k}$, k = 1, …, 13), the average (Avg), and consensus prediction (CoPr). Higher is better.
Perspectives Methods $A_{1}$ $A_{2}$ $A_{3}$ $A_{4}$ $A_{5}$ $A_{6}$ $A_{7}$ $A_{8}$ $A_{9}$ $A_{10}$ Avg CoPr
Happiness Base 0.80 0.12 0.27 0.38 0.56 0.35 0.44 0.44 0.31 0.55 0.42 0.45
D-LEMA 0.85 0.71 0.44 0.36 0.70 0.43 0.46 0.54 0.41 0.47 0.54 0.57
PADL 0.93 0.74 0.48 0.53 0.57 0.47 0.42 0.51 0.50 0.60 0.58 0.55
MaDL 0.91 0.77 0.44 0.38 0.70 0.47 0.46 0.54 0.48 0.47 0.56 0.58
Ours 0.94 0.80 0.54 0.55 0.69 0.51 0.53 0.54 0.52 0.64 0.63 0.62
Healthiness Base 0.77 0.19 0.14 0.47 0.87 0.36 0.43 0.44 0.51 0.54 0.47 0.56
D-LEMA 0.83 0.77 0.44 0.43 0.87 0.41 0.44 0.46 0.55 0.46 0.57 0.54
PADL 0.92 0.72 0.55 0.44 0.84 0.47 0.46 0.44 0.52 0.55 0.59 0.50
MaDL 0.89 0.77 0.44 0.44 0.90 0.41 0.44 0.46 0.55 0.46 0.58 0.55
Ours 0.92 0.75 0.56 0.52 0.90 0.49 0.48 0.54 0.64 0.61 0.64 0.58
Safety Base 0.58 0.65 0.36 0.36 0.61 0.37 0.56 0.40 0.32 0.53 0.47 0.51
D-LEMA 0.62 0.69 0.27 0.41 0.50 0.46 0.48 0.40 0.35 0.50 0.47 0.49
PADL 0.72 0.78 0.24 0.44 0.69 0.44 0.53 0.42 0.46 0.48 0.52 0.54
MaDL 0.63 0.63 0.27 0.32 0.61 0.38 0.46 0.42 0.36 0.52 0.46 0.56
Ours 0.72 0.80 0.38 0.48 0.71 0.54 0.53 0.50 0.52 0.58 0.58 0.61
Liveliness Base 0.79 0.60 0.53 0.46 0.76 0.30 0.35 0.36 0.53 0.57 0.53 0.55
D-LEMA 0.79 0.58 0.37 0.42 0.74 0.38 0.44 0.41 0.50 0.46 0.51 0.53
PADL 0.85 0.66 0.56 0.46 0.75 0.44 0.43 0.47 0.56 0.57 0.58 0.54
MaDL 0.78 0.56 0.35 0.40 0.76 0.34 0.48 0.42 0.47 0.47 0.50 0.56
Ours 0.87 0.68 0.57 0.53 0.80 0.49 0.48 0.51 0.62 0.61 0.62 0.59
Orderliness Base 0.50 0.64 0.39 0.45 0.86 0.31 0.39 0.34 0.31 0.49 0.47 0.52
D-LEMA 0.55 0.60 0.32 0.36 0.82 0.39 0.42 0.36 0.37 0.47 0.47 0.57
PADL 0.73 0.65 0.44 0.45 0.93 0.45 0.45 0.36 0.42 0.63 0.55 0.54
MaDL 0.61 0.60 0.34 0.36 0.86 0.37 0.47 0.36 0.37 0.49 0.48 0.58
Ours 0.74 0.71 0.52 0.55 0.94 0.47 0.54 0.44 0.56 0.62 0.61 0.62
Table 3. The accuracy metric is to evaluate results on the STREET dataset. We assess performance for individual annotator modeling (each annotator $A_{k}$, k = 1, …, 13), the average (Avg), and consensus prediction (CoPr). Higher is better.
Method S-Ha S-He S-Sa S-Li S-Or AMER
Full-PADL 0.58 0.59 0.52 0.58 0.55 0.79
Full-Ours 0.63 0.64 0.58 0.62 0.61 0.84
Sparse-PADL 0.43 0.42 0.38 0.41 0.40 0.58
Sparse-Ours 0.52 0.51 0.46 0.49 0.48 0.66
Table 4. Evaluation by accuracy for sparse scenarios (40% of annotations are randomly removed). S-Ha, S-He, S-Sa, S-Li, and S-Or represent five perspectives of STREET dataset: happiness, healthiness, safety, liveliness, and orderliness.
5.1. Implementation Details
Our image-specific model pipeline (Figure 2) uses ViT-G/14 from EVA-CLIP (Sun et al., 2023) as the encoder, with Image Q-Former initialized from InstructBLIP (Dai et al., 2023) (Frame Q-Former is the same in a video-specific pipeline from supplementary material, where Video Q-Former is initialized from Video-LLaMA (Zhang et al., 2023b)). Input images and video frames are resized to 224$\times$224 and normalized. The number of query tokens is set equal to the number of annotators, and each annotator’s classifier model uses an MLP. We train the model using the AdamW optimizer with an initial learning rate of 1e-4, a weight decay of 0.01, and gradient clipping with a maximum norm of 1.0. A linear warmup strategy is applied for the first 20% steps followed by cosine learning rate decay. The model is trained for up to 200 epochs with early stopping (patience = 25) to avoid overfitting. Training is conducted using distributed data parallelism (DDP) on four NVIDIA V100 GPUs.
Figure 3. A visualization analysis of the different image patches that annotators focused on in the STREET dataset. The annotators 4, 5, 6, 7, and 8 exhibit centralized focuses on a cute dog compared to other annotators.
Figure 4. A visualization analysis of the different video frames that annotators focused on in the AMER dataset. The annotator 1 exhibits focus on final frames (the frames contain different person), while annotators 4 and 13 focus on the middle frames.
5.2. Evaluation Metrics
To evaluate the performance of individual annotator modeling and consensus prediction (majority-votes the predictions of multi-annotators), we use accuracy (a standard metric in the multi-annotator learning) and $F_{1}$ score (Zhan et al., 2019) balancing precision and recall, suitable for potential class imbalance from uneven annotation densities, as in AMER (1,040 vs. 5,195 labels for annotators 1–10 vs. 11–13).
5.3. Quantitative Results
We analyze evaluation results on modeling individual annotators’ behavior patterns, their utility for consensus application, and applicability under sparse annotation scenarios.
Individual Annotator Modeling. Individual annotator modeling aims to capture the behavior patterns of different annotators. Results in Tables 2 and 3 show that our method consistently outperforms all baselines in both accuracy and $F_{1}$ score  [^2] See supplementary material for $F_{1}$ results on STREET dataset. across individual annotators on the STREET and AMER datasets. This validates the superiority of our approach in capturing individual annotator behavior patterns.
Consensus Application Benefits. Real-world applications often seek a single consensus label despite subjectivity among annotators. We evaluate consensus prediction to validate whether modeling individual annotators preserves valuable information for practical needs. As no definitive ground truth exists, we use majority vote over raw annotations as a proxy, acknowledging its potential biases. Existing baselines adopt different aggregation strategies: D-LEMA learns weighted fusion; PADL applies meta-learning; and MaDL jointly optimizes consensus and annotator classifiers. They may average annotator perspectives during training, potentially diminishing individual nuances. For a fair comparison, we apply unified majority voting over annotator-specific predictions from all methods rather than using their original aggregated outputs. Results (CoPr) in Tables 2, 3, and $F_{1}$ results 2 show that our method achieves superior consensus performance, suggesting that modeling individual annotators helps retain valuable information, potentially benefiting real-world consensus applications. Note: This experiment serves to validate practical utility of individual annotator modeling rather than to assert overall superiority.
Applicability under Sparse Annotations. To evaluate our model’s applicability under sparse annotation scenarios, we simulated real-world conditions by randomly removing annotations at various rates. As shown in Table 4, when 40% of annotations are removed [^3] See supplementary material for results of more sparse rates., our model’s average performance drops by 20.4%, whereas the best baseline PADL experiences a larger drop of 27.4%. Results suggest that our superiority stems from modeling inter-annotator correlations, which regularizes individual annotator representations, preventing overfitting to sparse labels and promoting consistency with shared patterns across annotators, to enhance robustness and generalization under sparse annotations.
5.4. Qualitative Results
We qualitatively analyze how the learned attention reflects annotators’ behavior patterns by visualizing cross-attention weights from Q-Former. These weights highlight the image patches or video frames that different annotators may focus on when making predictions.
As shown in Figure 3, on the STREET dataset (city impression classification), annotators exhibit distinct spatial focus: annotators 4, 5, 6, 7, and 8 focus more centrally on a dog in the image, while others do not. Correspondingly, these annotators also provided higher scores for the “happiness” dimension, suggesting that variations in focus on specific semantics may contribute to their differing judgments.
Figure 4 illustrates results on the AMER dataset (video emotion classification). Annotators differ in temporal focus: annotator 1 focuses on the final frames (45–50), while annotator 4 and 13 concentrate early frames (5–10). These patterns align with their predictions: “sad” for annotator 1, “worried” for annotator 4 and 13, suggesting that differences in the people or dialogues in the corresponding frame segments they focused on may underlie decision diversity.
Method S-Ha S-He S-Sa S-Li S-Or AMER
Base 0.42 0.47 0.47 0.53 0.47 0.28
w/ U-cls 0.50 0.56 0.49 0.56 0.49 0.61
w/o S-Attn 0.52 0.54 0.46 0.53 0.56 0.73
Pre-mv 0.58 0.57 0.56 0.58 0.60 0.43
Post-mv 0.62 0.58 0.61 0.59 0.62 0.60
Full (avg) 0.63 0.64 0.58 0.62 0.61 0.84
Table 5. Ablation study. The average performance (accuracy) of replacing modules, removing inter-annotator correlations, and a consensus prediction comparison with and without individual annotators’ behavior pattern modeling.
5.5. Ablation Study
We conduct an ablation study to validate our architectural design choices and assess the effect of individual behavior pattern modeling on consensus prediction (Table 5).
Architecture Choices. Removing Q-Former yields the Base model, and replacing individual classifiers with a unified classifier (w/ U-cls) both lead to significant performance drops, validating their effectiveness in our architecture.
Inter-Annotator Correlation Analysis. Disabling self-attention (w/o S-Attn) leads to clear performance degradation, underscoring the role of inter-annotator correlations as an implicit structural regularizer that improves generalization and robustness of individual annotator modeling.
Individual Behavior Modeling for Consensus. To investigate the role of individual behavior modeling in consensus prediction, we compare two strategies: (1) Pre-mv performs majority voting before modeling,i.e., applies global pooling over all Q-Former queries for a single prediction; (2) Post-mv first models each annotator individually and then aggregates their predictions. Post-mv consistently outperforms Pre-mv, especially on AMER. This demonstrates that modeling individual annotators’ behavior patterns preserves valuable information otherwise lost in early aggregation, potentially benefiting real-world consensus prediction or other practical applications.
6. Conclusion
This paper introduced a paradigm shift in multi-annotator learning from sample-wise aggregation to annotator-wise behavior modeling, and proposed a lightweight query-based architecture (QuMATL) to effectively model individual annotator behavior patterns. We contributed the STREET and AMER datasets with dense per-annotator labels for longitudinal annotator behavior understanding. Experiments demonstrate the superiority of our method in modeling individual annotators’ behavior patterns, their utility for consensus prediction, and applicability under sparse annotations. This work presents a novel view on recognizing and understanding the challenges in the multi-annotator learning field.
Figure 5. QuMATL architecture for video-specific pipeline. A frozen pre-trained encoder extracts frame features, which are compressed through the Frame Q-Former. With added frame position embedding, these interact with $1,\dots,n$ annotator-specific learnable queries in the Video Q-Former through cross-attention, producing annotator-specific features for classification. Different annotators’ cross-attention weights represent different annotator tendencies and provide visualization for analysis.
7. Supplementary Material
7.1. Overview
In this supplementary, we first illustrate the detailed video-specific architecture in Section 7.2. We provide the analysis of model efficiency in Section 7.3. Also, the extended discussion is provided in Section 7.4. Finally, we show the additional experimental results in Section 7.5.
7.2. Video-specific Architecture
Different from the image input pipeline in the main paper, Figure 5 presents the video input pipeline specifically designed for the AMER, etc. video datasets. Specifically, given a video input $V\in\mathbb{R}^{T\times H\times W\times 3}$, a frozen pre-trained image encoder (Sun et al., 2023) first extracts frame features, which are then further compressed by the image Q-Former using a specific number of compression queries, typically 32, to alleviate subsequent computational costs (Li et al., 2023). The frame position embedding is then added and input to the video Q-Former. Subsequently, we define the unique tendency for annotator $A_{k}$ ($k=1,\dots,n$) through each learnable query in a Q-Former, then all annotator-specific queries in cross-attention simultaneously interact with input features to capture their diverse tendencies and obtain annotator-specific feature. Finally, the specific annotator features are mapped through a fully connected layer to an appropriate feature dimension and then connected to each annotator’s corresponding classifier to output classification results.
For visualization, the focus points to video input are based on all frames of the video. Different annotators’ varying levels of focus on different frames indicate their tendency differences: annotator 1 shows more focus on the final frames, while annotator n focuses more on the middle frames. This dataset is video emotion classification. Analyzing the annotation and visualization results, we see that annotators labeled emotions “anger” and “worried”. The differences in the frame segments they focused on might partially contribute to the variations in their judgments.
Table 6. Model efficiency analysis. In this table, we report the number of parameters and the average inference time per sample. Lower values for both metrics indicate higher model efficiency.
Models Parameters (M) Average Processing Time (s)
D-LEMA $214.18$ $5.64$
PADL $168.94$ $4.59$
MaDL $201.37$ $5.06$
Ours $\bf{106.02}$ $\bf{4.28}$
7.3. Model Efficiency Analysis
This section evaluates the model efficiency from two aspects: model complexity and inference time. For model complexity, we use the number of parameters as the evaluation metric; for inference time, we compute the average processing time per sample (Zhang et al., 2023c, d, 2022; Zhang, 2024; Sha et al., 2024). To ensure a fair comparison, Table 6 compares the performance of different methods in the image input pipeline (see Figure 2 in the main paper). Specifically, we evaluate the efficiency of different methods on one perspective of the STREET dataset with 10 annotators. It is worth noting that different methods use different backbone networks. To ensure fairness, we standardize the backbone network to ResNet-34 for all methods. In Table 6, we observe that our model has fewer parameters than the other models, while also maintaining competitive average processing time. This validates our claim in the paper that, compared to the baselines that create separate conventional models for each annotator, our architecture demonstrates significant advantages in model efficiency.
7.4. Discussion
We supplement some extended discussions about our work here. Our architecture design superiorly balances effectiveness and complexity, provides an accompanying visualization analysis of annotator tendencies through cross-attention weights. This has potential value and interest for understanding different annotator judgments as described in the main paper. Currently, it is introduced as an accompanying function rather than a main contribution of our paper, but in the future, we hope it can be developed into a mature interpretability solution. We plan to explore to enhance it, which may need pixel-level annotations indicating specific regions that annotators focus on during the annotation process, although this could be expensive. We could also explore quantitative evaluation ways to this further interpretability, such as feature importance ranking consistency, attention-based fidelity metrics, and human intuition consistency assessments through user studies, etc, to further enhance its value and contribution to the community.
In the ablation study of our main paper, we validated an insight that modeling multiple annotator tendencies can obtain more annotator information and demonstrated its effectiveness through experimental results based on consensus aggregation using majority voting. In future work, we plan to extend more similar experimental scenarios and validations to enhance the explanation of how multi-annotator tendency learning helps and adds benefits to specific applications (Zhang et al., 2014).
7.5. Additional Results
We provide additional experimental results about visualization analysis of annotator tendencies on both AMER and STREET datasets.
As shown in Figure 6, on the AMER dataset, annotators demonstrate different tendencies through varying focus on different frames. Sample 1 reveals the differences: annotators 5, 10, and 11 show similarly higher focus on the final frames (45-54), while annotator 12 focuses more on middle frames (30-35). They all predict the correct label “happy”, but the different focus positions indicate that annotator 12’s preference pattern for happiness in sample 1 differs from most annotators. Sample 2 shows: annotators 2, 8, and 13 exhibit similarly higher focus on the final frames (45-55), while annotator 12 focuses more on early frames (5-10). They all predict “sad” but annotator 12 demonstrates a different preference pattern.
As shown in Figure 7, on the STREET dataset, annotators exhibit different tendencies through varying focus on different semantic elements within the same input image. In sample 1 from the orderliness perspective, the focus regions show differences: annotator 5 focuses more intensively on graffiti in the image compared with other annotators. The results show that prediction and ground truth are highly consistent in the emotion reflected in the semantics, corresponding to this annotator’s expected lowest score. In sample 2 of the healthiness perspective, most annotators focus on uncleaned garbage and leaves, and only annotator 6 shows more focus on the surrounding environment, like cars and buildings, which might influence the result of not giving a low score.
(a) Sample 1.
(b) Sample 2.
Figure 6. Additional sample experimental results visualize the annotator tendencies on the AMER video dataset. The tendencies of multi-annotators reveal distinct preferences.
(a) Sample 1 for orderliness perspective.
(b) Sample 2 for healthiness perspective.
Figure 7. Additional sample experimental results visualize the annotator tendencies on the STREET image dataset. The tendencies of 10 annotators reveal distinct preferences.